# On the nature of the spin glass transition Gesualdo Delfino SISSA – Via Bonomea 265, 34136 Trieste, Italy

### v1

@@ 4 Conclusion
Recently, the first exact access to the spin glass regime in finite dimension has been obtained in $d=2$ [7]. This was achieved exploiting conformal invariance in the scattering framework to implement the replica method in an exact way and to obtain the renormalization group fixed points of the Ising spin glass. The problem involves $N$ copies (real replicas) of the system whose correlations (overlaps) characterize spin glass physics and are measured in the simulations, as well as $n$ auxiliary replicas (for each copy) needed to implement quenched disorder within the replica method. Very remarkably, the exact solution revealed the mechanism through which the necessary $N$-independence emerges precisely in the limit $n\to 0$ implied by the replica method. The presence of a line of fixed points, while unexpected because rarely realized in physically relevant contexts, immediately proved able to explain features of Ising spin glass criticality observed in numerical studies in $d=2$.
In this paper we observed that the existence of a line of fixed points in $d=2$ reveals the enhancement to a one-generator continuous internal symmetry (shift symmetry). We noted similarities and differences with the enhancement from discrete to continuous symmetry known for some antiferromagnets with infinitely many degenerate ground states. While the line of fixed points implies a condition of exact marginality which does not persist in $d>2$, the continuous symmetry that it reveals characterizes the model in all dimensions. In $d=2$ the symmetry enhancement explains the absence, so far mysterious, of a finite temperature transition to a spin glass phase. In $d>2$, instead, the spontaneous breaking of the continuous symmetry is allowed and yields a spin glass phase with an overlap order parameter which, for fixed temperature and disorder strength, takes continuous values in an interval. An order parameter with such a feature characterizes Parisi mean field solution of the model with infinite-range interactions, which corresponds to $d=\infty$. We argued that, in a subtle way, $N$-independence at $n=0$ allows the mean field order parameter – which involves only the auxiliary replicas – to inherit the characteristic feature of the order parameter for the real replicas.

### v2

@@ 4 Conclusion
Recently, the first exact access to the spin glass regime in finite dimension has been obtained in $d=2$ [7]. This was achieved exploiting conformal invariance in the scattering framework to implement the replica method in an exact way and to obtain the renormalization group fixed points of the Ising spin glass. The problem involves $N$ copies (real replicas) of the system whose correlations (overlaps) characterize spin glass physics and are measured in the simulations, as well as $n$ auxiliary replicas (for each copy) needed to implement quenched disorder within the replica method. Very remarkably, the exact solution revealed the mechanism through which the necessary $N$-independence emerges precisely in the limit $n\to 0$ implied by the replica method. The presence of a line of fixed points, while unexpected because rarely realized in physically relevant contexts, immediately explains properties of Ising spin glass criticality observed in numerical studies in $d=2$, namely the presence and specific features of nonuniversality, and the almost-free-bosonic behavior of [69].
In this paper we observed that the existence of a line of fixed points in $d=2$ reveals the enhancement to a one-generator continuous internal symmetry (shift symmetry). We noted similarities and differences with the enhancement from discrete to continuous symmetry known for some antiferromagnets with infinitely many degenerate ground states. While the line of fixed points implies a condition of exact marginality which does not persist in $d>2$, the continuous symmetry that it reveals characterizes the model in all dimensions. In $d=2$ the symmetry enhancement immediately explains the absence of a finite temperature transition to a spin glass phase observed in the numerical simulations. In $d>2$, instead, the spontaneous breaking of the continuous symmetry is allowed and yields a spin glass phase with an overlap order parameter which, for fixed temperature and disorder strength, takes continuous values in an interval. An order parameter with such a feature characterizes Parisi mean field solution of the model with infinite-range interactions, which corresponds to $d=\infty$. We argued that, in a subtle way, $N$-independence at $n=0$ may allow the mean field order parameter – which involves only the auxiliary replicas – to inherit the characteristic feature of the order parameter for the real replicas.